Represent the sum of first terms of the series 3+33+333+.........

using the sigma notation.



1
Expert's answer
2022-03-22T04:15:31-0400
"S= 3(1+11+111+1111+...n^{th} term)."

Let the nth term be,

"tn=1+10+100+...10^{n-1}"

This is a G.P. with first term 1 and common ratio of 10.

Thus, the sum of n terms,

"S= 3\sum_{i=1}^{n} t_{n}=3\sum_{i=1}^{n}\frac{10^n-1}{9}"

Hence,

"S= 3\sum\frac{10^n}{9}-3\sum\frac{1}{9}"

"S= 3[\frac{10(1-10^n)}{(1-10)\times9}-(\frac{1}{9})\times n"

"= 3[10\times\frac{10^n-1}{81}-\frac{n}{9}]"

Solving this we get,

"S= [10\times \frac{10^n-1}{27}-\frac{n}{3}]"


Hence, the sum of first n terms of the series is,


"= \frac{10^{n+1}-10-9n}{27}"